# SimLTD: Simple Supervised and Semi-Supervised Long-Tailed Object Detection

### v1

@@ 4.5 Ablation Experiments
Design of SimLTD Baseline The design of SimLTD is centered on an intuitive multi-stage training strategy combined with standard data augmentations, without bells and whistles, to overcome the class imbalances in both head and tail datasets. Table 4 shows the contributions of RFS and Copy-Paste in establishing a more robust baseline than was previously possible in the existing literature, by using the simple Faster R-CNN detector with ResNet-50 FPN. We also showcase the viability of multi-stage learning over the naïve single-stage training procedure on the whole long-tailed dataset, which yields markedly worse results.
@@
How Many Shots to Sample for ? Figure 8 illustrates the impact on mAP and AP as a function of sampled shots for fine-tuning with , ranging from 10 to “all” meaning the entire long-tailed training set. The aim of this experiment is to optimize for accuracy on rare classes while mitigating catastrophic forgetting on pre-trained head representations. We analyze the “knee in the curve” and find that 30-shots balance the trade-off between the two metrics. Figure 3 visualizes that with 30-shots, we sample the entire tail distribution containing 20 or fewer instances per class and include a mixture of head categories for exemplar replay. Moving to the left of this “sweet spot” with -shots, we observe marked reductions in mAP, indicating adverse forgetting on head classes from insufficient samples. Moving to the right of the sweet spot, we see a precipitous drop in AP in response to the overwhelming amount of head samples.
@@ 5 Conclusion
We introduced SimLTD, a simple and versatile framework for supervised and semi-supervised long-tailed detection. Standing out from existing work, SimLTD delivers excellent performance and scalability by achieving new state-of-the-art results on the challenging LVIS v1 benchmark, without requiring auxiliary image-level supervision. We hope the practitioner finds utility in our multi-stage training approach and for our work to spur future research aimed at pushing the performance envelope of long-tailed detection.

### v2

@@ 4.5 Ablation Experiments
Design of SimLTD Baseline
The design of SimLTD is centered on an intuitive multi-stage training strategy combined with standard data augmentations, without bells and whistles, to overcome the class imbalances in both head and tail datasets. Table 4 shows the contributions of RFS and Copy-Paste in establishing a more robust baseline than was previously possible in the existing literature, by using the simple Faster R-CNN detector with ResNet-50 FPN. We also showcase the viability of multi-stage learning over the naïve single-stage training procedure on the whole long-tailed dataset, which yields markedly worse results.
@@
How Many Shots to Sample for ? Figure 8 illustrates the impact on mAP and AP as a function of sampled shots for fine-tuning with , ranging from 10 to “All” meaning the entire long-tailed training set. The aim of this experiment is to optimize for accuracy on rare classes while mitigating catastrophic forgetting on pre-trained head representations. We analyze the “knee in the curve” and find that 30-shots balance the trade-off between the two metrics. Figure 3 visualizes that with 30-shots, we sample the whole tail distribution containing 20 or fewer instances per class and include a mixture of head categories for exemplar replay. Moving to the left of this “sweet spot” with -shots, we observe marked reductions in mAP, indicating adverse forgetting on head classes from insufficient samples. Moving to the right of the sweet spot, we see a precipitous drop in AP in response to the overwhelming amount of head samples.
@@ 5 Conclusion
We introduced SimLTD, a simple and versatile framework for supervised and semi-supervised long-tailed detection. Standing out from existing work, SimLTD delivers excellent performance and scalability by achieving new state-of-the-art results on the challenging LVIS v1 benchmark—without requiring auxiliary image-level supervision. We hope the practitioner finds utility in our multi-stage training approach and for our work to spur future research aimed at pushing the performance envelope of long-tailed detection.